It takes 51.0 J to raise the temperature of an 8.80 g piece of unknown metal from 13.0°C to 24.8°C what is the specific heat for the metal?
Answer #1

Hi,

Hope you are doing well.


__________________________________________________________________________

Specific heat capacity is defined by the amount of heat required to raise the temperature of a 1 kg substance through 1oC.

Given that, Total heat required to raise the temperature, H=51\;J

Change in temperature, \Delta T=T_{2}-T_{1}=24.8^{o}C-13^{o}C=11.8^{o}C

Mass of the substance, m=8.80\;g=8.80\times10^{-3}\;kg

(1 g = 10-3 kg)

Therefore, Specific heat capacity of the metal is given by,

C=\frac{H}{m\Delta T}=\frac{51\;J}{8.80\times10^{-3}\;kg\times11.8^{o}C}

\boldsymbol{\therefore C=491.14\;J/\left (kg^{o}C \right )}


__________________________________________________________________________

Hope this helped for your studies. Keep learning. Have a good day.

Feel free to clear any doubts at the comment section.


Please don't forget to give a thumps up.

Thank you. :)